OpenStudy (anonymous):

solve by factoring and using the principle of zero products. 1.) 3x^2-27x+60=0 2.) x(x-1)=12

OpenStudy (anonymous):

1) (3x - 12)(x -5) = 0 x = 5, 4

OpenStudy (anonymous):

2)(x - 4)(x + 3) = 0 x = 4, -3

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

would be easier to factor if you wrote \[3x^2-27x+60=0\] \[3(x-9+20)=0\] first